# How Does a Deep Neural Network Look at Lexical Stress?

### v1

@@ I Introduction
While deep learning based architectures have transformed many aspects of speech processing, those models often exhibit an intransparent prediction process, obscuring how specific decisions are made. This work explores the application of one technique for shedding light on such architectures through the lens of a well-studied phenomenon in speech science and speech processing: lexical stress. Lexical stress refers to the prominence, accentuation, or emphasis of a particular syllable within a word. In many languages, lexical stress is crucial for communication as it can can serve as the primary basis for a contrast in meaning. For example, in English, the words “PROtest” (noun; stress on initial syllable) and “proTEST” (verb; stress on final) have contrasting stress patterns but the same phonemes (with predictable changes to the quality of the initial vowel depending on stress).
@@
A long tradition of phonetics research has examined the (relative) temporal, spectral, and energetic properties of stressed and unstressed syllables [5, 6]; see Cutler [4] and Gay [7] for reviews. Building on this work, automatic stress detection methods have used acoustic features such as fundamental frequency, amplitude, and energy [19, 29]. Bayesian classifiers with multivariate Gaussian distributions have been applied to enhance stress detection accuracy, utilizing acoustic features such as peak-to-peak amplitude, normalized energy, and vowel duration [30, 31]. Similar to other areas of speech processing, recent work using deep learning architectures has significantly improved over these more traditional approaches. Convolutional Neural Networks (CNNs) and Multi-Distribution Deep Neural Networks (MD-DNNs) outperformed traditional models in both English and Arabic stress detection in second language speech [25, 18]. Additionally, an attention-based neural network combined with data augmentation through a neural network based text-to-speech system has been introduced to detect lexical stress errors in second-language English learners, achieving significant improvements in precision and recall [15].
@@
Although these advancements in deep learning have led to significant improvements in speech processing, they also introduce challenges, particularly in understanding classifier decisions and gaining insight into how these models operate. This issue of interpretability is crucial. Better understanding the acoustic features that drive model predictions can help us verify that the model will be reliable when applied to novel data (which may have different distributions for critical features). Neural network interpretability has broadly become an active research area, with a number of techniques being explored across different domains. Here, we focus on Layerwise Relevance Propagation (LRP) [2], a technique that was developed to provide insights into the decision-making processes of these complex models by understanding what specific portions of the input influence the model’s output. We use this technique to investigate how CNNs – which have demonstrated superior performance in lexical stress detection – derive their predictions from the spectral features in their input.
@@
This paper offers several novel contributions. First, a new speech dataset was automatically gathered without human annotation, using a pipeline built around deep learning systems. Second, we present a state-of-the-art lexical stress classifier and show its ability to generalize to novel words, including words from previously unseen languages. Lastly, we employ LRP to analyze the spectral features these classifiers use to achieve high accuracy.
@@ II Dataset Construction
Previous studies on lexical stress often depended on stress labels assigned by human annotators [25, 15, 29] or were assigned based on the canonical stress for the orthographic word taken from pronunciation dictionaries such as the CMU dictionary [18]. However, manual annotation is costly and can introduce inter-annotator inconsistency, and the use of a single citation form can yield errors for orthographically ambiguous minimal pair words (e.g., protest). Prior work also failed to explicitly document measures to prevent lexical overlap between training and evaluation splits, leaving open the possibility that identical word types were seen during both stages. Reported accuracy may therefore overestimate true generalization. To avoid these concerns, we created a novel dataset, using a fully automatic pipeline, without utilizing human annotations.
@@ II.1 Selection of English Disyllabic Words
We prompted ChatGPT 4o [10] to generate 30 pairs of disyllabic words that form stress minimal pairs: pairs of words with the same spelling but different meaning depending on the primary stress placement. Four of these items could not be incorporated into other parts of our pipeline (the part-of-speech tagger, discussed below, failed to correctly annotate these forms). The remaining 26 pairs are address, conduct, content, contrast, convert, decrease, digest, entrance, export, extract, import, increase, insult, object, perfect, present, progress, project, protest, rebel, record, refuse, reject, subject, suspect, and transfer; the initial and final syllable stress variants of these words will be referred to as “minimal pairs.” We also asked ChatGPT 4o to generate a larger set of 250 disyllabic words that do not have stress minimal pairs (e.g., initial stress WALLet, final stress exTEND). After excluding one error, we retained 249 words (124 initial stress, 125 final stress). We refer to this as the set of “no minimal pair words.”
@@ II.2 Extraction of Words from Existing Corpora
These words were then extracted from three existing English speech corpora. The first dataset is LibriSpeech [22], a widely-used corpus consisting of approximately 1,000 hours of read speech derived from audiobooks. The second dataset is The Supreme Court corpus [28], which includes recordings of oral arguments made before the United States Supreme Court. This dataset provides speech data from a highly formal and legal context, offering variation in speaker formality and legal terminology that can influence acoustics associated with lexical stress. The third dataset is TED-LIUM [9], derived from TED Talks. It includes over 400 hours of speech by speakers from various language backgrounds and nationalities, discussing a wide range of topics. The diversity in speaking styles and contexts in TED-LIUM further enriches the phonetic variation in our dataset.
@@
The processing pipeline is shown in Figure 1. First, we used the Montreal Forced Aligner [21] to align the transcript with the audio, extracting precise time intervals for each target word in our minimal pairs and no minimal pair words lists in each of the datasets. After identifying the location of target minimal pair words within the audio, we applied the SpaCy [1] part-of-speech tagging system to the entire audio in order to determine whether each word was used as noun or verb. This classification follows linguistic conventions for English disyllabic words, where nouns typically carry primary stress on the initial syllable, referred to as initial stress (IS), and verbs carry primary stress on the final syllable, referred to as final stress (FS) [19]. For the no minimal pair, we used annotations from the CMU Pronouncing Dictionary [18], as the stress placement is unambiguous in this case.
Words were extracted precisely at the forced aligner start time and lasted 0.5 seconds. This fixed width was necessary for the use of the CNN architecture; we verified that this window was sufficiently large to encompass both syllables of every disyllabic word. We also split each minimal pair word into syllables using an automatically syllabified version of the CMU Pronouncing dictionary  [^1] The syllabified dictionary and syllabifier can be accessed at http://sourceforge.net/p/p2tk/. and phoneme-level time stamps from the forced aligner.
@@
Overall, we had 6,659 samples (124 types) and 3,095 samples (125 types) for initial and final stress in the no minimal pairs set, respectively, and 5,547 samples (40 types) and 1,260 samples (40 types) for initial and final stress in minimal pairs, respectively.
@@ II.3 Phonetic Properties of Extracted Samples
To provide an initial verification that our procedures yielded sets of disyllabic words that contrasted in stress location, two well-known correlates of lexical stress were examined: amplitude (vowels in stressed syllables are louder than those in unstressed syllables) and duration (vowels in stressed stressed syllables are longer than those in unstressed syllables) [5, 6]. For each sample, we calculated the ratio of the properties of the vowels of the initial versus final syllables. These relative measures allow us to control for potential differences in speech rate and amplitude across samples. Our samples reliably showed the expected differences. Samples with initial stress had high ratios of amplitude (mean 0.6, s.d. 0.18) and duration (mean 0.53, s.d. 0.15), reflecting relatively loud, long vowels in stressed initial versus unstressed final syllables. In contrast, samples with final stress showed lower mean ratios for amplitude (mean 0.29, s.d. 0.14) and duration (mean 0.35, s.d. 0.12), reflecting relatively quiet, short vowels in unstressed initial versus stressed final syllables. To confirm that this is a reliable difference between sample types, we used bootstrap resampling (with 1,000 replicates) to estimate the 95% confidence interval for differences in mean ratios between initial versus final stress samples. These differences were significant (amplitude mean difference: 0.305; 95% CI (0.302,0.309); p<p<.0001; duration mean difference: 0.178, 95% CI (0.175,0.18), $p<$ .0001). This provides independent confirmation that the samples in this data set have contrasting stress patterns.
@@ III.1 Model Architectures
In our study, we investigated different CNN architectures to develop a binary classifier for identifying the initial or final stress of stress-minimal word pairs. The input to all models consists of magnitude spectrograms obtained using the Short-Time Fourier Transform (STFT), following z-score normalization. The STFT was computed with a sampling rate of 16,000 Hz, using a Hamming window of size 0.02 seconds and a window stride of 0.01 seconds. We experimented with several well-known CNN architectures. The LeNet-5 model consists of 3 connected convolutional layers with average pooling between the layers, followed by 2 dense layers in a 2x2 configuration [17]. The VGG11, VGG16, and VGG19 architectures consist of 8, 13, and 16 convolutional layers, respectively, with max pooling between the layers, followed by an additional 3 dense layers [27]. Finally, the ResNet18 model consists of 18 convolutional layers with residual connections to address the vanishing gradient problem during training [8].
Figure 2: Spectrograms and corresponding attributions for the words “PREsent” and ‘preSENT” using the VGG16 model. Orange vertical lines represent phoneme boundaries (the label on each line denotes the segment to the left of the boundary; 1 and 0 denote stressed and unstressed vowels, respectively). The blue vertical line represents the end of the initial syllable. Panel (a) shows the spectrogram of “PREsent” (IS), while panel (e) shows the spectrogram of ”preSENT” (FS). Panels (b)-(d) and (f)-(h) display attributions using various methods: $\mathbf{LRP_{\epsilon}},\mathbf{LRP_{\alpha 1}}\text{ and }\mathbf{LRP_{CMP}}$ respectively, shown over the spectrograms. These LRP variants effectively capture initial versus final stress contrasts, highlighting features at different scales. The $\mathbf{LRP_{\alpha 1}}$ (panel (c) vs. (g))and $\mathbf{LRP_{CMP}}$ methods (panel (d) vs. (h)) show clearly different patterns depending on the location of stress.
III.2 Data Splitting and Training Process
@@
The training set was constructed using 201 out of the 249 word types from the no minimal pair words. Data augmentation techniques were employed to enlarge this set of examples. For each sample, we added a low-pass filtered version of the sample (cutoff frequency at 3000 Hz) as well as three versions of the sample mixed with noise. A sample of multi-talker babble from the VOiCES Corpus [23]) was combined with the original sample at three different levels (20 dB, 10 dB, and 3 dB SNR). These augmentations aimed to simulate real-world conditions, enhancing the robustness of our models. After the augmentation, we had 29,810 and 13,510 samples from the initial and final stress, respectively, corresponding to the 201 no minimal pair word types. The augmented dataset will be referred to as the “no minimal pairs Train.”
The remaining 48 no minimal pair word types were divided into two lists of 24 words each, with one list used for the validation (i.e., hyper-parameter tuning;“no minimal pairs Validation”) and the other for the test set (“no minimal pairs Test”). The 26 minimal pairs words will be referred to as “minimal pairs Test”. Note that two test sets were used (no minimal pairs Test and minimal pairs Test) as the minimal pairs set may contain errors; stress was determined solely by part-of-speech tagging, which may be inaccurate. In contrast, the tagging for the no minimal pair words is more reliable; as the word form is unambiguous, the lexically-determined stress pattern can be accurately identified.
The models were trained using standard procedures for CNNs. To stabilize training, we employed an Adam Optimizer [12] with a learning rate schedule and batch normalization. To address class imbalance (i.e., more initial versus final stress examples), we employed focal loss [20], modifying the standard cross-entropy loss to focus more on hard-to-classify examples (the gamma parameter was tuned to 2 based on validation set performance).
Figure 3: Further illustrative examples of spectrograms and corresponding attributions for the words “subject,” “record” and “refuse” with different primary stress location, using the VGG16 model. Vertical lines and annotation follow Figure 2. Panel (a)-(d) show spectrograms and corresponding $\mathbf{LRP_{CMP}}$ heatmaps for the word ”subject” with initial and final stress, respectively. Similarly, panels (e)-(l) display spectrograms and heatmaps of the words “record” and “refuse,” with different stress types.

### v2

@@ I Introduction
While deep learning based architectures have transformed many aspects of speech processing, those models often exhibit an nontransparent prediction process, obscuring how specific decisions are made. This work explores the application of one technique for shedding light on such architectures through the lens of a well-studied phenomenon in speech science and speech processing: lexical stress. Lexical stress refers to the prominence, accentuation, or emphasis of a particular syllable within a word. In many languages, lexical stress is crucial for communication as it can serve as the primary basis for a contrast in meaning. For example, in English, the words “PROtest” (noun; stress on initial syllable) and “proTEST” (verb; stress on final) have contrasting stress patterns but the same phonemes (with predictable changes to the quality of the initial vowel depending on stress).
@@
A long tradition of phonetics research has examined the (relative) temporal, spectral, and energetic properties of stressed and unstressed syllables [fry1955duration, fry1958experiments]; for reviews, see cutler, gay1978physiological, gordon2017acoustic, and vanHeuvenStress. Building on this work, automatic stress detection methods have used acoustic features such as fundamental frequency, amplitude, and energy [lieberman1960some, 1415269]. Bayesian classifiers with multivariate Gaussian distributions have been applied to enhance stress detection accuracy, utilizing acoustic features such as peak-to-peak amplitude, normalized energy, and vowel duration [1168788, 607932]. Similar to other areas of speech processing, recent work using deep learning architectures has significantly improved over these more traditional approaches. Convolutional Neural Networks (CNNs) and Multi-Distribution Deep Neural Networks (MD-DNNs) outperformed traditional models in both English and Arabic stress detection in second language speech [shahin2016automatic, LI201828]. An attention-based neural network combined with data augmentation through a neural network based text-to-speech system has been introduced to detect lexical stress errors in second-language English learners, achieving significant improvements in precision and recall [korzekwa2020detection]. Classifiers applied to self-supervised learning representations from wav2vec 2.0 rather than traditional acoustic features also show high performance [bentum2024processing].
@@
Although these advancements in deep learning have led to significant improvements in speech processing, they also introduce challenges, particularly in understanding classifier decisions and gaining insight into how these models operate. This issue of interpretability is crucial. Better understanding the acoustic features that drive model predictions can help us verify that the model will be reliable when applied to novel data (which may have different distributions for critical features). Neural network interpretability has broadly become an active research area, with a number of techniques being explored across different domains. Here, we focus on Layerwise Relevance Propagation (LRP) [bach2015pixel], a technique that was developed to provide insights into the decision-making processes of these complex models by understanding what specific portions of the input influence the model’s output. We use this technique to investigate how CNNs – which have demonstrated superior performance in lexical stress detection – derive their predictions from the spectral features in their input.
@@
This paper offers several novel contributions. First, a new speech dataset was automatically gathered without human annotation, using a pipeline built around deep learning systems. Second, we present a state-of-the-art model for lexical stress classification in disyllabic words and show its ability to generalize to novel words. Lastly, we employ LRP to analyze the spectral features these classifiers use to achieve high accuracy.
@@ II Dataset Construction
Previous studies on lexical stress often depended on stress labels assigned by human annotators [shahin2016automatic, korzekwa2020detection, 1415269] or were assigned based on the canonical stress for the orthographic word taken from pronunciation dictionaries such as the CMU dictionary [LI201828]. However, manual annotation is costly and can introduce inter-annotator inconsistency, and the use of a single citation form can yield errors for orthographically ambiguous minimal pair words (e.g., protest). Prior work also failed to explicitly document measures to prevent lexical overlap between training and evaluation splits, leaving open the possibility that identical word types were seen during both stages [shahin2016automatic, korzekwa2020detection, 1415269, LI201828]. Reported accuracy may therefore overestimate true generalization. To avoid these concerns, we created a novel dataset, using a fully automatic pipeline, without utilizing human annotations [^1] The complete dataset, including all audio samples, is publicly available at https://huggingface.co/datasets/MLSpeech/lexical_stress_dataset..
@@ II.1 Selection of English Disyllabic Words
We prompted ChatGPT 4o [hurst2024gpt] to generate 30 pairs of disyllabic words that form stress minimal pairs: pairs of words with the same spelling but different meaning depending on the primary stress placement. Five of these items could not be incorporated into other parts of our pipeline (the part-of-speech tagger, discussed below, failed to correctly annotate these forms). The remaining 25 pairs are address, conduct, content, contrast, convert, decrease, digest, export, extract, import, increase, insult, object, perfect, present, progress, project, protest, rebel, record, refuse, reject, subject, suspect, and transfer; the initial and final syllable stress variants of these words will be referred to as “minimal pairs.” We also asked ChatGPT 4o to generate a larger set of 250 disyllabic words that do not have stress minimal pairs (e.g., initial stress WALLet, final stress exTEND). After excluding one error, we retained 249 words (124 initial stress, 125 final stress). We refer to this as the set of “no minimal pair words.”
@@ II.2 Extraction of Words from Existing Corpora
These words were then extracted from three existing English speech corpora. The first dataset is LibriSpeech [7178964], a widely-used corpus consisting of approximately 1,000 hours of read speech derived from audiobooks. The second dataset is The Supreme Court corpus [spaeth2014supreme], which includes recordings of oral arguments made before the United States Supreme Court. This dataset provides speech data from a highly formal and legal context, offering variation in speaker formality and legal terminology that can influence acoustics associated with lexical stress. The third dataset is TED-LIUM [hernandez2018ted], derived from TED Talks. It includes over 400 hours of speech by speakers from various language backgrounds and nationalities, discussing a wide range of topics. The diversity in speaking styles and contexts in TED-LIUM further enriches the phonetic variation in our dataset.
The processing pipeline is shown in Figure 1. First, we used a forced alignment algorithm to align the transcript with the audio. Specifically we used Montreal Forced Aligner (MFA) [mcauliffe2017montreal] to align LibriSpeech and Supreme Court. MFA works at time-resolution of 10 millisecond and shown by rousso2024tradition to outperform recent end-to-end ASR-based aligners such as WhisperX [bain2023whisperx] and MMS [pratap2024scaling]. For TED-LIUM, which already includes word-level timestamps, we used the provided alignments. Importantly, unlike MFA-aligned corpora (LibriSpeech and Supreme Court), where word boundaries are adjacent (e.g., “hello 0–0.3 second”, “my 0.3–0.4”), the TED-LIUM timestamps may be non-adjacent (e.g., “hello 0–0.3”, “my 0.35–0.4”), leaving short pauses between words.
Once the start time of each word is obtained, a 0.5-second window is opened. This window size was chosen to ensure that both syllables of every disyllabic word were fully included and to maintain a uniform input length compatible with our CNN architecture. We retained only those whose total duration was less than 0.5 second and zero-padded them as needed to reach exactly 0.5 second. Each minimal pair word was further segmented into syllables using a syllabification algorithm [^2] The syllabified dictionary and syllabifier can be accessed at http://sourceforge.net/p/p2tk/.. In addition, phoneme-level timestamps derived from the forced aligner were incorporated.
@@
After identifying the location of target minimal pair words within the audio, we applied part-of-speech tagging system [SpaCy; Annex2020] to the entire audio in order to determine whether each word was used as noun or verb. The tagging follows linguistic conventions for English disyllabic words, where nouns typically carry primary stress on the initial syllable, referred to as initial stress (IS), and verbs carry primary stress on the final syllable, referred to as final stress (FS) [lieberman1960some]. For the no minimal pair, we used annotations from the CMU Pronouncing Dictionary [LI201828], as the stress placement is unambiguous in this case. During this process, we observed that certain words, specifically address, content, export, extract, and decrease, were particularly affected by inaccuracies in the part-of-speech tagging system. Many of these errors involved incorrectly labeling words with initial stress instead of final stress. To correct these inconsistencies, we manually reviewed and relabeled all instances of these words.
@@
Overall, we had 7,446 samples (124 types) and 3,263 samples (125 types) for initial and final stress, respectively, in the no minimal pairs set, and 5,476 samples and 1,715 samples for initial and final stress, respectively, in the minimal pairs set, consisting of a total of 25 pairs.
@@ II.3 Phonetic Properties of Extracted Samples
To provide an initial verification that our procedures yielded sets of disyllabic words that contrasted in stress location, two well-known correlates of lexical stress were examined: amplitude (vowels in stressed syllables are louder than those in unstressed syllables) and duration (vowels in stressed stressed syllables are longer than those in unstressed syllables) [fry1955duration, fry1958experiments]. For each sample, we calculated the ratio of the properties of the vowels of the initial versus final syllables. These relative measures allow us to control for potential differences in speech rate and amplitude across samples. Our samples reliably showed the expected differences. Samples with initial stress had high ratios of amplitude (mean 0.6, s.d. 0.18) and duration (mean 0.53, s.d. 0.15), reflecting relatively loud, long vowels in stressed initial versus unstressed final syllables. In contrast, samples with final stress showed lower mean ratios for amplitude (mean 0.29, s.d. 0.14) and duration (mean 0.35, s.d. 0.12), reflecting relatively quiet, short vowels in unstressed initial versus stressed final syllables. To confirm that this is a reliable difference between sample types, we used bootstrap resampling (with 1,000 replicates) to estimate the 95% confidence interval for differences in mean ratios between initial versus final stress samples. These differences were significant (amplitude mean difference: 0.305; 95% CI (0.302,0.309); p<p<.0001; duration mean difference: 0.178, 95% CI (0.175,0.18), $p<$ .0001). This provides independent confirmation that the samples in this data set have contrasting stress patterns.
Figure 2: Spectrograms and corresponding attributions for the words ”PREsent” and ”preSENT” using the VGG16 model. Orange vertical lines represent phoneme boundaries, with each label centered within its phoneme segment; 1 and 0/2 denote stressed and unstressed vowels, respectively. The blue vertical line represents the end of the initial syllable. Panel (a) shows the spectrogram of ”PREsent” (IS), while panel (e) shows the spectrogram of ”preSENT” (FS). Panels (b)-(d) and (f)-(h) display attributions using various methods: $\mathbf{LRP_{\epsilon}},\mathbf{LRP_{\alpha 1}}\text{ and }\mathbf{LRP_{CMP}}$ respectively, shown over the spectrograms. These LRP variants effectively capture initial versus final stress contrasts, highlighting features at different scales. The $\mathbf{LRP_{\alpha 1}}$ (panel (c) vs. (g))and $\mathbf{LRP_{CMP}}$ methods (panel (d) vs. (h)) show clearly different patterns depending on the location of stress. Note that the phoneme labels differ slightly across the two rows; these reflect natural differences in pronunciation between initial and final stress members of this word pair (i.e., stress-related changes to vowel quality).
@@ III.1 Model Architectures
In our study, we investigated different CNN architectures to develop a binary classifier for identifying the initial or final stress of stress-minimal word pairs. The input to all models consists of magnitude spectrograms obtained using the Short-Time Fourier Transform (STFT), following z-score normalization (see Figures 2 and 3 for examples). The STFT was computed with a sampling rate of 16,000 Hz, using a Hamming window of size 0.02 seconds and a window stride of 0.01 seconds. We experimented with several well-known CNN architectures. The LeNet-5 model consists of 3 connected convolutional layers with average pooling between the layers, followed by 2 dense layers in a 2x2 configuration [726791]. The VGG11, VGG16, and VGG19 architectures consist of 8, 13, and 16 convolutional layers, respectively, with max pooling between the layers, followed by an additional 3 dense layers [simonyan2014very]. Finally, the ResNet18 model consists of 18 convolutional layers with residual connections to address the vanishing gradient problem during training [He2016CVPR].
III.2 Data Splitting
@@
The training set was constructed using 201 out of the 249 word types from the no minimal pair words. Data augmentation techniques were employed to enlarge this set of examples. For each sample, we added a low-pass filtered version of the sample (cutoff frequency at 3000 Hz) as well as three versions of the sample mixed with noise. A sample of multi-talker babble from the VOiCES Corpus [richey2018voices] was combined with the original sample at three different levels (20 dB, 10 dB, and 3 dB SNR). These augmentations aimed to simulate real-world conditions, enhancing the robustness of our models. After the augmentation, we had 29,830 and 13,510 samples from the initial and final stress, respectively, corresponding to the 201 no minimal pair word types. The augmented dataset will be referred to as the “no minimal pairs train.”
The remaining 48 no minimal pair word types were divided into two lists of 24 words each, with one list used for the validation (i.e., hyper-parameter tuning; “no minimal pairs validation”) and the other for the test set (“no minimal pairs test”). The 25 minimal pairs words will be referred to as “minimal pairs test”. Note that two test sets were used (“no minimal pairs test” and “minimal pairs test”) as the minimal pairs set may contain errors; stress was determined solely by part-of-speech tagging, which may be inaccurate. In contrast, the tagging for the no minimal pair words is more reliable; as the word form is unambiguous, the lexically-determined stress pattern can be accurately identified. See Table 1 for dataset composition.
Table 1: Composition of the train (before and after data augmentations), validation, and test sets. Each set consists of disjoint word types (no overlap across splits). The table reports the number of unique word types and the total number of samples corresponding to initial stress (IS) and final stress (FS) words.
Set #Word types #IS samples (w. aug. ) #FS samples (w. aug. )
no minimal pairs train 201 5966 (29,830) 2702 (13,510)
no minimal pairs validation 24 763 260
no minimal pairs test 24 717 301
minimal pairs test 25 5476 1715
III.3 Training Process
The models were trained using standard procedures for CNNs. To stabilize training, we employed an Adam Optimizer [kingma2014adam] with a learning rate $5\times 10^{-4}$ and batch normalization. To address class imbalance (i.e., more initial versus final stress examples), we employed the focal loss function [lin2017focal], modifying the standard cross-entropy loss to focus more on hard-to-classify examples. Focal loss introduces a modulating factor $(1-p_{t})^{\gamma}$ to the cross-entropy objective, reducing the relative loss contribution of well-classified examples and amplifying that of misclassified ones. Formally, it is defined as:
$FL(p_{t})=-(1-p_{t})^{\gamma}\log(p_{t})~,$ (1)
where $p_{t}$ denotes the predicted probability for the true class, and $\gamma\geq 0$ is the focusing parameter that controls the rate at which easy examples are down-weighted. In our experiments, we set $\gamma=2$ based on validation performance, which provided a good balance between stability and discrimination.
Each model was trained for up to 50 epochs, with early stopping triggered after 15 epochs without validation improvement. All experiments were conducted on 8 NVIDIA A40 GPUs.